OpenStudy (anonymous):

Given that 2^(2x+3) x 7^x-2 = 2^3x (49^x), evaluate 14^x

ganeshie8 (ganeshie8):

start wid dividing the whole equation wid \(2^{2x+3} 7^{x-2}\)

OpenStudy (anonymous):

done it

ganeshie8 (ganeshie8):

\(\large 2^{2x+3} 7^{x-2} = 2^{3x} (49^x)\) \(\large 1 = \frac{2^{3x} (49^x)}{2^{2x+3} 7^{x-2}}\)

ganeshie8 (ganeshie8):

use exponent properties, and get rid of denominator in right hand side

ganeshie8 (ganeshie8):

also see that you can write 49 as 7^2

OpenStudy (anonymous):

seen it i'm ready to move on

ganeshie8 (ganeshie8):

\(\large 2^{2x+3} 7^{x-2} = 2^{3x} (49^x)\) \(\large 1 = \frac{2^{3x} (49^x)}{2^{2x+3} 7^{x-2}}\) \(\large 1 = \frac{2^{3x} (7^{2x})}{2^{2x+3} 7^{x-2}}\) \(\large 1 = \frac{2^{x} (7^{x}) 7^2}{2^{3} }\) \(\large 1 = \frac{14^{x} 7^2}{2^{3} }\)

OpenStudy (anonymous):

Thanks for teaching me...

ganeshie8 (ganeshie8):

np :)